OpenStudy (anonymous):

combining like terms simplify each expression do not solve the problem lol -5n+3(6+7n) = 2n+9n is this correct?

OpenStudy (anonymous):

LHS = -5n + 18 + 21n = 16n + 18 first step is remove the brackets by multiplying each term in the brackets by + 3

OpenStudy (agreene):

-5n + 18 + 21n (21-5)n + 18 16n+18 factoring, if you want it simplified: 2(8n+9)

OpenStudy (anonymous):

Jimmy are you solving this problem?

OpenStudy (anonymous):

16n + 18 = 2(8n + 9)

OpenStudy (anonymous):

no just simplifying the left hand part

OpenStudy (anonymous):

when solving leave LHS as 16n + 18